If a person invests $260 at 4% annual interest, find the approximate value of the investment at the end of 5 years?
Answer by ikleyn(52906)

.
If a person invests $260 at 4% annual interest, find the approximate value of the investment at the end of 5 years?
~~~~~~~~~~~~~~~~~


Assuming simple interest, the value of the investment at the end of 5 years is


260*(1+0.04*5) = 260*(1 + 0.2) = 260*1.2 = 312 dollars.